# DCF-DS: Deep Cascade Fusion of Diarization and Separation for Speech Recognition under Realistic Single-Channel Conditions

### v1

@@ IV-E Performance Comparison on LibriCSS Dataset
Table IV presents the performance of the proposed methods on the single-channel LibriCSS dataset. It is important to note that we test the pre-trained model directly without fine-tuning for domain adaptation on the LibriCSS dataset. This is relatively challenging since the LibriCSS dataset includes overlapping speech data with overlap ratios ranging from 0 to 40, and it contains utterances with long periods of silence (0L), which are almost unseen in our training set. The results show that DCF-DS still achieves strong performance on the LibriCSS dataset, with trends consistent with those observed in Table II. Specifically, using longer window lengths, MIMO-SE module, and re-clustering all contribute to improvements, demonstrating the good generalization ability of the DCF-DS method. Finally, we achieve a tcpWER of 4.43 on the single-channel LibriCSS dataset. From Tables II and IV, it is shown that although the number of output nodes in DCF-DS is limited to 3 or 4, the method still performs well in multi-speaker scenarios (e.g., 4-8 speakers in NOTSOFAR-1 eval set and 8 speakers in LibriCSS dataset). This also indicates that DCF-DS effectively mitigates the SDCI issue.
@@
Table V presents the cpWER results of representative methods on the LibriCSS dataset under the single-channel constraint. The first three rows show the recognition results using different conventional clustering methods, including agglomerative hierarchical clustering (AHC) [30], variational Bayesian hidden Markov model with x-vectors (VBx) [35], and spectral clustering (SC) [52]. It can be observed that, without separation, the recognition performance is generally poor (cpWER over 30). The Transcribe-to-Diarize [68] method uses an ASR model to generate timestamps, achieving a cpWER of 11.6. The system in [25], which combines TS-VAD and speech extraction, achieves a cpWER of 18.8. The TS-SEP method reaches cpWERs of 11.61 and 7.87 using E2E and E2E-SSL as ASR models, respectively. CSS-AD [26] achieves a cpWER of 6.2 based on the E2E-SSL ASR model. Finally, our proposed DCF-DS achieves cpWER of 4.43 with the E2E-SSL ASR model, demonstrating the state-of-the-art recognition performance on the single-channel LibriCSS dataset.

### v2

@@ IV-E Performance Comparison on LibriCSS Dataset
Table IV presents the performance of the proposed methods on the single-channel LibriCSS dataset. It is important to note that we test the pre-trained model directly on the LibriCSS dataset without domain adaptation. This is relatively challenging since the LibriCSS dataset includes recordings with overlap ratios ranging from 0 to 40. It also contains utterances with long periods of silence (0L subset), which are almost unseen in our training set. The results show that DCF-DS still achieves strong performance on the LibriCSS dataset, with trends consistent with those observed in Table II. Specifically, longer window lengths, MIMO-SE module, and re-clustering all contribute to improvements, demonstrating the good generalization ability of the DCF-DS method. Finally, we achieve a cpWER of 4.43 on the single-channel LibriCSS dataset. From Tables II and IV, it is shown that although the number of output nodes in DCF-DS is limited to 3 or 4, the method still performs well in multi-speaker scenarios (e.g., 4-8 speakers in NOTSOFAR-1 eval set and 8 speakers in LibriCSS dataset). This also indicates that DCF-DS effectively mitigates the SDCI issue.
@@
Table V presents the cpWER results of representative methods on the LibriCSS dataset under the single-channel constraint. The first three rows show the recognition results using different conventional clustering methods as the front-end, including agglomerative hierarchical clustering (AHC) [30], variational Bayesian hidden Markov model with x-vectors (VBx) [35], and spectral clustering (SC) [52]. It can be observed that, without separation, the recognition performance is generally poor (cpWER over 30). The Transcribe-to-Diarize [69] method uses an ASR model to generate timestamps, achieving a cpWER of 11.6. The system in [25], which combines TS-VAD and speech extraction, achieves a cpWER of 18.8. The TS-SEP method reaches cpWERs of 11.61 and 7.87 using E2E and E2E-SSL as ASR models, respectively. CSS-AD [26] achieves a cpWER of 6.2 based on the E2E-SSL ASR model. Finally, our proposed DCF-DS achieves cpWER of 4.43 with the E2E-SSL ASR model, demonstrating the state-of-the-art recognition performance on the single-channel LibriCSS dataset.